X-ray radiation falls in the wavelength region of 1.00×10^-11 to 1.00×10^-8 meters.
What is the wavelength of X-ray radiation that has an energy of 1.99×10^-17 kJ/photon?
Wavelength =_______m
1
Expert's answer
2020-05-13T14:21:43-0400

E=hv.

v=E/h.

lambda=c/v.

1.99 x 10-7/6.62 x 10-34 = 3,0060423e26 1/s.

lambda = 3 x 10^ 8 / 3,0060423e26=

1e-18 km. =1e-15 m.